OpenStudy (anonymous):

y=25x^2+10x-24. Find the zeros of this equation.

OpenStudy (anonymous):

Factor, you get (5x + 6)(5x - 4) = 0 Set each factor equal to zero, you get 5x + 6 = 0, or x = -6/5 5x - 4 = 0, you get x = 4/5 Two solutions for x.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

welcome.